Do the domain of the function f defined by √(3-x) /(x-2) is R - {2}

1
Expert's answer
2021-11-05T06:05:37-0400

The domain of the function is the set of the all values of the argument where the function exists.

In the given case: "f(x)={\frac{\sqrt{3-x}} {x-2}}"

The function exists when "3-x\u2265 0" and "(x-2)\not=0"

So, the domain is "x\in (-\infty;3]"\ {2}, not R \ {2}